callielovesyhuu:

A rock is rolled down a 300-foot steep incline. After 1 second, the rock is traveling 26 feet per second. After 2 seconds, the rock is traveling 52 feet per second. Assume the relationship between time, t, and speed, s, is linear. Write an equation in slope-intercept form describing this relationship. Use ordered pairs of the form (time, speed).

callielovesyhuu:

I need help

Hanna:

y =32x+12 i think

TEXECUTIONER:

(1,44) and (4,140) 140 - 44


= 96/3 = 32 = m 4 - 1 y = mx + b 44 = 32(1) + b 12 = b y = 32x + 12 <==ANSWER

TEXECUTIONER:

that's if you didn't know how to do it:)

callielovesyhuu:

thank you <3

TEXECUTIONER:

No problem;)

callielovesyhuu:

fan for fan?

TEXECUTIONER:

Sure!

callielovesyhuu:

im already ur fan <3

TEXECUTIONER:

haha thanks ;)

callielovesyhuu:

welcome